OpenStudy (anonymous):

If a number is increased by 50% then the resultant number is decreased by 40%. What was the intial number if the final number is 8 less than the original.

OpenStudy (phi):

If a number is increased by 50% translated "a number" to x (x for unknown) increased by 50% means add 50% of x to x: x + 50% of x 50% of x is written 0.50*x or just 0.5x so a number is increased by 50% becomes: x+0.5x if we factor out the x: (1+0.5)x add 1+0.5 to get 1.5, and we get 1.5x a number is increased by 50% becomes 1.5x

OpenStudy (phi):

then the resultant number is decreased by 40%. if we have a number (call it y for the time being) and decrease it by 40% we do this: y - 40% of y or y - 0.4*y or (1-0.4)y or 0.6y that means to decrease a number by 40% , multiply it by 0.6 then the resultant number is decreased by 40%. the "resultant number" from the first step is 1.5x to decrease it by 40% we multiply by 0.6 0.6*1.5x = 0.9x

OpenStudy (phi):

What was the intial number if the final number is 8 less than the original. the initial number is x the final number is 0.9x the final number is 8 less than the original. 8 less than the original means original - 8 or x - 8 final number is 8 less than the original 0.9x = x - 8 or 0.9x = x -8 can you solve for x ?

OpenStudy (anonymous):

Yes. I can solve for x.. thanks.